OpenStudy (anonymous):

Please help ASAP XD The ages of members of a seniors curling club are normally distributed, with a mean of 63 years and a standard deviation of 4 years. What percent of the curlers is in each of the following age groups? A) Between 55 and 63 years old B) Between 67 and 75 years old C) Older than 75 years

OpenStudy (kirbykirby):

You will need to standardize the values. Like, for a) , say X = age Then we are finding \(P(55 < X <63)\). By standardizing, we find:\[P\left(\frac{55-63}{4}<\frac{X-63}{4}<\frac{63-63}{4}\right)=P\left(-2<Z<0\right)\]

OpenStudy (anonymous):

OKay thanks, i think i know what im doing now XD

OpenStudy (kirbykirby):

Ok hehe good then. b) is similar to a). c), you are just finding P(X > 75), and standardize that again